Find the y and x intercepts of the equation
6y-12=66x

Answer by algebrahouse.com(1659)

Find the y and x intercepts of the equation
6y - 12 = 66x

To find the x-intercept, put 0 in for y:
6(0) - 12 = 66x {substituted 0 in for y}
-12 = 66x {multiplied 6 by 0}
x = -12/66 {divided both sides by 66}
x = -2/11 {reduced}
x-intercept = -2/11
coordinates are (-2/11,0)

To find the y-intercept, put 0 in for x:
6y - 12 = 66(0) {substituted 0 in for x}
6y - 12 = 0 {multiplied 66 by 0}
6y = 12 {added 12 to both sides}
y = 2 {divided both sides by 6}
y-intercept = 2
coordinates are (0,2)
www.algebrahouse.com